OpenStudy (afloridagirl):

Factor x2 − 2x + 3. (x − 3)(x − 1) (x + 3)(x + 1) (x − 3)(x + 1) Prime

OpenStudy (afloridagirl):

i believe its c or a

OpenStudy (mjdennis):

Or try multiplying your answers, and see which one works!

jhonyy9 (jhonyy9):

x^2 -2x +3 = 0 so because the discriminant smaller then zero D= 4 -12 = -8 so what mean this ?

OpenStudy (afloridagirl):

im sorry idk @jhonyy9

jhonyy9 (jhonyy9):

so when the discriminant is smaller than zero so this mean that this quadratic have not real solutions ok. ?

OpenStudy (afloridagirl):

im sorry i had to leave thank you for trying to explain everything to me almost positive that i am just to dumb to understand basic stuff @jhonyy9